OpenStudy (micahm):

Evaluate the function f(x) = 2x − 2, when x = −1 -4 -3 0 4 I think its A

OpenStudy (dangerousjesse):

Plug in \((2\times -1) -2\)

OpenStudy (dangerousjesse):

2 times -1= -2 so 2-2=?....

OpenStudy (dangerousjesse):

-2-2 i mean

OpenStudy (micahm):

so could you break it down step by stept I suck at math

OpenStudy (dangerousjesse):

You plug -1 in \(F(x)=2x-2:\) \(F(-1)=2\times -1 -2\) Using \(\small\color{red}{\textsf{PEMDAS}}\), Solve. \(2\times -1= -2\) \(-2-2=-4\) \(F(x)=-4\)

OpenStudy (micahm):

ok thanks